# Constraint-Aware Reinforcement Learning via Adaptive Action Scaling

### v1

@@ Abstract
I Introduction
Reinforcement Learning (RL) has demonstrated remarkable success across a range of domains, including Atari games [1], robotics [2, 3], and long-horizon strategy games [4, 5]. This success is significantly facilitated by exploratory behavior, which allows agents to discover effective behaviors. However, such exploratory behaviors often lead to the violation of constraints imposed on the controlled system. While constraint violations are tolerable in simulated environments and games where resets are free, they pose serious risks in real-world applications [6]. Violating safety constraints can lead to irreversible damage or system failure. To address this issue, Safe Reinforcement Learning (Safe RL) [7] has emerged as a critical area of research that aims to minimize constraint violations during both training and deployment.
Safe RL methods can be broadly categorized into two groups: safe exploration and constrained RL. Safe exploration techniques aim to prevent the agent from taking actions that violate safety constraints. These methods typically rely on prior knowledge of the system dynamics and feasible safe states to construct control barrier functions [8, 9, 10], or model predictive shields [11, 12, 13]. Although effective, their applicability is limited by the need for detailed prior information about the system dynamics, an assumption that often does not hold in early learning stages or tasks where system dynamics are unknown.
Figure 1: Overview of cost-aware action scaling. The RL agent proposes an action that would result in the center of mass (COM) exceeding the velocity threshold (left). The regulator (blue) intervenes by scaling the action, keeping the velocity of the COM within the safe zone while allowing progress on the task. The yellow circles highlight the velocity threshold for the COM, illustrating how the regulator enforces a safety constraint while preserving task performance.
Constrained RL instead allows the agent to learn both reward and cost signals online, without requiring knowledge of the system dynamics. The agent is trained to maximize cumulative rewards while minimizing constraint violations. Common approaches include Lagrangian-based methods that solve the dual formulation of the constrained optimization problem [14, 15, 16, 17], and budget-based methods maintain a running estimate of remaining safety allowance, adjusting the agent’s behavior to respect cost limits over time [18, 19]. However, a core limitation of these methods is the difficulty of balancing reward and cost within a single policy. Conflicting gradients can cause the agent to behave either too conservatively or unsafely, leading to instability, constraint violations, or poor performance [17, 20].
@@ I Introduction
We evaluate our approach on several dynamical systems from Safety Gymnasium [21] and the Safety-Critical environments from [22]. Our method achieves the highest Return-to-Cost (RC) ratio [23], reducing constraint violations by up to 126 times over recent safe RL baselines [17, 19, 24, 25, 26]. In summary, our contributions are:
  • We propose a modular safe RL framework that decouples reward maximization and safety enforcement via a cost-aware regulator that scales actions based on predicted violations.
  • Our model-free approach integrates seamlessly into standard off-policy RL pipelines such as SAC [27] and TD3 [28], improving safety without compromising exploration.
  • We achieve state-of-the-art performance on safety benchmarks, with up to 126 times fewer constraint violations and the highest RC ratios across tasks.
II Related Work
@@
Safe RL aims to enable agents to maximize task rewards while minimizing constraint violations. Existing approaches broadly fall into two categories: safe exploration methods, which intervene externally to prevent unsafe actions during training, and constrained RL methods, which embed cost objectives directly into policy optimization.
@@
Safe Exploration Methods. Safe exploration techniques intervene during training to prevent agents from entering unsafe states. Early methods, such as [29], employed uncertainty modeling through Gaussian Processes to restrict exploration and ensure safety during optimization. Later approaches introduced safety layers [30, 31] and predictive safety filters [11, 12, 13, 22] that anticipate and block risky actions based on pre-trained layers or model predictive control (MPC). Control Barrier Function (CBF)-based strategies [8, 9, 10] encode safety constraints directly through differentiable control barriers to guarantee that the agent’s actions remain within certified safe sets throughout training. [32] extend shielding methods to continuous domains by leveraging approximate dynamics models, enabling probabilistic safety guarantees during exploration. [33] leveraged model-based RL and offline collected data to develop reachability-based safety layers to ensure safe actions for navigation scenarios. [23, 34] assumes access to an offline dataset for pretraining a cost critic along with a recovery policy, which is then fixed during online learning, limiting its applicability in settings where collecting sufficient offline data is challenging or costly.
While our method shares some conceptual similarity with these safe exploration approaches—modifying actions to maintain safety—it differs fundamentally by relying on online-learned cost predictions rather than external models or handcrafted safe sets, and by smoothly scaling actions instead of hard blocking or overwriting them, preserving the agent’s exploratory behavior.
@@
Constrained RL Approaches. Constrained RL methods integrate cost minimization into policy learning itself. Lagrangian-based algorithms [14, 16, 17] optimize dual formulation balancing rewards and costs, while budgeted RL approaches [18, 19] include the remaining cost budget in the state representation, allowing the agent to adapt its behavior based on how much cost it can still afford. Risk-sensitive formulations, such as CVaR-CPO [35], enforce safety by constraining the conditional value-at-risk of cumulative costs, ensuring attention to costly violations. Reachability-based methods like RESPO [25] estimate the probability of reaching safe regions and optimize policies to persistently satisfy constraints or recover when outside the feasible set. Constraint-Conditioned Policy Optimization [36] enables zero-shot generalization to unseen cost thresholds by conditioning the policy and value functions on constraint levels using a variational inference objective. Bi-level optimization frameworks such as SRCPO [26] address the nonlinearity of risk measures by optimizing over dual variables, achieving strong constraint satisfaction in continuous control tasks. Safety Editor [24] instead trains two separate Soft Actor-Critic (SAC) [27] agents: a utility maximizer and a safety editor that modifies unsafe actions, allowing it to fully overwrite the original action when necessary.
Compared to these methods, our approach offers a lightweight, modular alternative: instead of embedding constraints into the policy loss, relying on delicate dual updates, or training a second actor to overwrite unsafe actions, we regulate actions externally using learned cost estimators. This continuous scaling preserves exploration while enabling seamless integration into standard off-policy RL pipelines.
@@ III Preliminaries
Constrained Reinforcement Learning. Constrained RL refers to the problem of maximizing task rewards while satisfying explicit constraints on agent behavior. The standard formalism is through Constrained Markov Decision Processes (CMDPs) [37], where constraints are expressed using cost functions separate from the reward.
Constrained Markov Decision Processes. A CMDP augments the MDP with a cost function $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}_{\geq 0}$ that quantifies safety violations. The objective is to maximize return while keeping the expected cumulative cost below a budget $\chi$:
$\max_{\pi}\;$ $\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot|s)}[Q^{\pi}(s,a)]$
s.t. $\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot|s)}[Q_{c}^{\pi}(s,a)]\leq\chi,$ (CMDP)
where the cost value functions are
$V_{c}^{\pi}(s)$ $=\mathbb{E}_{\pi}\Big[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\,\Big|\,s_{0}=s\Big],$ (1)
$Q_{c}^{\pi}(s,a)$ $=\mathbb{E}_{\pi}\Big[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\,\Big|\,s_{0}=s,\,a_{0}=a\Big].$ (2)
Cost Budget. The budget $\chi$ specifies the maximum allowable expected cumulative cost and is typically treated as a human-selected threshold that reflects task-specific safety requirements [17]. In this work, we assume a stricter setting by eliminating the cost budget, i.e., setting $\chi=0$, similar to [25]. This corresponds to a hard-safety regime that aims to achieve minimal constraint violations during learning.
@@
$\pi^{*}$ $=\arg\max_{\pi}\;\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot|s)}[Q^{\pi}(s,a)]$ (3)
s.t. $\;\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot|s)}[Q_{c}^{\pi}(s,a)]=0.$ (4)
@@ IV Methodology
IV-A Split Architecture for Reward and Cost Optimization
@@
IV-B Action Modulation via Regulator Scaling
@@
where $\odot$ denotes element-wise multiplication, where each component of the action vector is multiplied by a scaling factor between 0 (high risk, large attenuation) and 1 (safe, no attenuation), smoothly reducing potentially unsafe actions proportional to predicted risk. This element-wise modulation attenuates each component of the action based on its risk profile, reducing the magnitude of high-risk components. Unlike hard safety constraints that may override agent behavior, this approach preserves the agent’s exploration behavior and allows stable off-policy learning. In many robotic domains, safety costs naturally increase with the magnitude of control inputs: high torques in manipulators accelerate wear and overheating, and large contact forces in legged robots risk joint damage. By designing costs that capture this structure, practitioners can align the regulator with system-specific safety considerations, making scaling an intuitive and broadly applicable mechanism for enforcing safety.
IV-C Learning Objectives and Updates
@@
$Q_{r}(s_{t},\tilde{a}_{t})\;\leftarrow\;$ $\;r(s_{t},\tilde{a}_{t})$
$+\gamma\,\mathbb{E}_{\begin{subarray}{c}s_{t+1}\sim p(\cdot|s_{t},\tilde{a}_{t})\\ a_{t+1}\sim\pi(\cdot|s_{t+1})\end{subarray}}\!\left[Q_{r}(s_{t+1},\tilde{a}_{t+1})\right],$ (6)
@@
ensuring that policy learning reflects the actual dynamics induced by the regulated action $\tilde{a}_{t}$. Our framework is algorithm-agnostic and can be integrated with any off-policy actor-critic method. For entropy-regularized algorithms such as SAC, the corresponding entropy term may be included in the actor objective. In our experiments, we demonstrate compatibility with both SAC and Twin Delayed DDPG (TD3)[28].
Cost Learning. The cost critic is also trained on the scaled actions using a TD-style Bellman backup [38]:
$Q_{c}(s_{t},\tilde{a}_{t})\;\leftarrow\;$ $\;c(s_{t},\tilde{a}_{t})$
$+\gamma\,\mathbb{E}_{\begin{subarray}{c}s_{t+1}\sim p(\cdot|s_{t},\tilde{a}_{t})\\ a_{t+1}\sim\pi(\cdot|s_{t+1})\end{subarray}}\!\left[Q_{c}(s_{t+1},\tilde{a}_{t+1})\right],$ (8)
@@
$\mathcal{L}_{\text{reg}}=\;\mathbb{E}_{s_{t}\sim\mathcal{D},\,a_{t}\sim\pi(\cdot|s_{t})}\Big[$ $\beta\cdot Q_{c}(s_{t},\tilde{a}_{t})$ (9)
$-\lambda\cdot\log\rho_{\theta}(s_{t},a_{t},\hat{c}_{t})\,\Big].$
@@
Algorithm 1 Cost-Aware Action Scaling (training loop)
1:Environment $\mathcal{E}$, replay buffer $\mathcal{D}$, actor $\pi_{\phi}$, regulator $\rho_{\theta}$, reward critic $Q_{r}$, cost critic $Q_{c}$
2:Initialize network parameters and target networks
3:for each interaction step do
4:  Observe $s$ and sample $a\sim\pi_{\phi}(\cdot\mid s)$
5:  $\hat{c}\leftarrow\max\!\big(Q_{c}^{1}(s,a),\,Q_{c}^{2}(s,a)\big)$ $\triangleright$ predicted cost
[…]
7:  Execute $\tilde{a}$, obtain $(r,c,s^{\prime})$, store $(s,\tilde{a},r,c,s^{\prime})$ in $\mathcal{D}$
8:end for
9:for each gradient step do
10:  Sample minibatch from $\mathcal{D}$
11:  Update $Q_{r}$ and $\pi_{\phi}$ via Eqs. (6)–(7)
12:  Update $Q_{c}$ (Eq. 8)
13:  Update $\rho_{\theta}$ (Eq. 9)
14:  Polyak-average target networks
15:end for
@@
Gradient Flow and Modularity. To ensure clean modularity, we detach the scaling weights $\rho_{\theta}(s_{t},a_{t},\hat{c}_{t})$ from the computational graph when updating both the reward and cost critics, preventing gradients from flowing through the regulator. Similarly, the actor receives no gradients from the regulator, learning purely from task returns. Moreover, the regulator is updated independently via its own objective, ensuring that reward maximization and safety modulation remain decoupled. The full training procedure is summarized in Algorithm 1.
@@
Implementation Details. The proposed method integrates a standard off-policy RL agent with a lightweight action regulator network for constraint satisfaction. The RL agent follows its baseline implementation without modification. The regulator is a feedforward neural network with two hidden layers of sizes [256, 256] and ReLU activations, outputting element-wise scaling factors between 0 and 1 via a sigmoid activation to modulate the agent’s actions. The regulator is trained using predicted costs from twin cost critics. Each cost critic is a feedforward network with two hidden layers of sizes [256, 256] and Tanh activations, taking state-action pairs as input and outputting a scalar cost prediction.
V Experiments
Figure 3: Performance comparison on the Safety Gymnasium locomotion environments. Each method is averaged over three independent runs; bold lines indicate the mean, and shaded areas show the standard deviation. Our methods (SAC-REG and TD3-REG) consistently achieve the best trade-off between return and cumulative constraint cost across all environments. Top: Episode return. Middle: Cumulative safety cost. Bottom: Return-to-Log-Cost ratio. Our methods outperform strong baselines, including PPO-PID [17], SIMMER [19], SRCPO [26], RESPO [25], and SEDITOR [24]. SIMMER is omitted from the Swimmer plot as it consistently yields negative returns, moving opposite to the target velocity.

### v2

@@ Abstract
1 Introduction
Figure 1: Overview of cost-aware action scaling. The RL agent proposes an action that would result in the center of mass (COM) exceeding the velocity threshold (left). The regulator (blue) scales the action, keeping the velocity of the COM within the safe zone while allowing progress on the task. The yellow circles highlight the velocity threshold for the COM.
Reinforcement Learning (RL) has demonstrated remarkable success across a range of domains, including Atari games [mnih2015human], robotics [gu2017deep, wu2023daydreamer], and long-horizon strategy games [silver2017mastering, vinyals2019grandmaster]. This success is significantly facilitated by exploratory behavior, which allows agents to discover effective behaviors. However, such exploratory behaviors often lead to the violation of system constraints. While such violations are tolerable in simulation and games with free resets, they pose serious risks in real-world applications [amodei2016concrete]. Violating safety constraints can lead to irreversible damage or system failure. To address this issue, Safe Reinforcement Learning (Safe RL) [garcia2015comprehensive] aims to minimize constraint violations during both training and deployment.
Safe RL methods can be broadly categorized into two groups: safe exploration and constrained RL. Safe exploration techniques aim to prevent the agent from taking actions that violate safety constraints. These methods typically rely on prior knowledge of the system dynamics and feasible safe states to construct control barrier functions [ames2019control, dai2023safe, zhang2023spatial], or model predictive shields [banerjee2024dynamic, dawood2025l4dc, agha2024exploring]. Although effective, their applicability is limited by the need for detailed prior information about the system dynamics, an assumption that often does not hold in early learning stages or tasks where system dynamics are unknown.
Constrained RL allows the agent to learn both reward and cost signals online, without requiring knowledge of the system dynamics. The agent is trained to maximize cumulative rewards while minimizing constraint violations. Common approaches include Lagrangian-based methods [achiam2017constrained, tessler2018reward, ray2019benchmarking, stooke2020responsive], and budget-based methods [sootla2022saute, sootla2022enhancing]. However, a core limitation of these methods is the difficulty of balancing reward and cost within a single policy. Conflicting gradients can cause the agent to behave either too conservatively or unsafely, leading to instability, constraint violations, or poor performance [stooke2020responsive, navon2022multi].
@@ 1 Introduction
We evaluate our approach on several dynamical systems from Safety Gymnasium [ji2023safety]. Our method achieves the highest Return-to-Cost (RC) ratio [thananjeyan2021recovery], reducing constraint violations by up to 126 times over recent safe RL baselines [stooke2020responsive, sootla2022enhancing, yu2022towards, ganai2023iterative, kim2024spectral]. In summary, our contributions are: (i) We propose a modular safe RL framework that decouples reward maximization and safety enforcement via a cost-aware regulator that scales actions based on predicted violations. (ii) Our model-free approach integrates seamlessly into standard off-policy RL pipelines such as SAC [haarnoja2018soft] and TD3 [fujimoto2018addressing], improving safety without compromising exploration. (iii) We achieve state-of-the-art performance on safety benchmarks, with up to 126 times fewer constraint violations and the highest RC ratios across tasks.
2 Related Work
@@
Existing approaches fall into two categories: safe exploration methods, which prevent unsafe actions, and constrained RL methods, which embed cost objectives directly into policy optimization.
@@
Safe Exploration Methods. Early methods, [sui2015safe], employed uncertainty modeling through Gaussian Processes to restrict exploration and ensure safety. Later approaches introduced safety layers [dalal2018safe, sheebaelhamd2021safe] and predictive safety filters [banerjee2024dynamic, dawood2025ral, agha2024exploring, tian2024reinforcement] that prevent risky actions based on pre-trained layers or model predictive control (MPC). Control Barrier Function (CBF)-based strategies [ames2019control, dai2023safe, zhang2023spatial] employ differentiable barriers to keep actions within certified safe sets. [goodall2024leveraging] extend shielding methods to continuous domains by leveraging approximate dynamics models, enabling probabilistic safety guarantees during exploration. [selim2022safe] leveraged model-based RL and offline collected data to develop reachability-based safety layers to ensure safe actions for navigation scenarios. [thananjeyan2021recovery, zhang2023safe] assumes access to an offline dataset for pretraining a cost critic along with a recovery policy, which is then fixed during online learning, limiting its applicability in settings where collecting sufficient offline data is challenging or costly. While our method also modifies actions to maintain safety, it differs fundamentally by relying on online-learned cost predictions rather than external models or handcrafted safe sets, and by smoothly scaling actions instead of hard blocking or overwriting them, preserving the agent’s exploratory behavior.
@@
Constrained RL Approaches. Lagrangian-based algorithms [achiam2017constrained, ray2019benchmarking, stooke2020responsive] optimize dual formulation balancing rewards and costs, while budgeted RL [sootla2022saute, sootla2022enhancing] include the remaining cost budget in the state representation, allowing the agent to adapt its behavior based on how much cost it can afford. Risk-sensitive formulations, such as CVaR-CPO [zhang2024cvar], constrain the conditional value-at-risk of cumulative costs, ensuring attention to costly violations. Reachability-based methods like RESPO [ganai2023iterative] estimate the probability of reaching safe regions and optimize policies to satisfy constraints or recover when outside the feasible set. Bi-level optimization frameworks such as SRCPO [kim2024spectral] address the nonlinearity of risk measures by optimizing over dual variables, achieving strong constraint satisfaction in continuous control tasks. Safety Editor [yu2022towards] trains two Soft Actor-Critic (SAC) [haarnoja2018soft] agents: a utility maximizer and a safety editor that modifies unsafe actions, allowing it to fully overwrite the original action when necessary. Compared to these methods, our approach offers a lightweight, modular alternative: instead of embedding constraints into the policy loss, relying on delicate dual updates, or training a second actor to overwrite unsafe actions, we regulate actions externally using learned cost critics. This continuous scaling preserves exploration while enabling seamless integration into standard off-policy RL pipelines.
@@ 3 Preliminaries
Constrained Markov Decision Processes. A CMDP [altman2021constrained] augments the MDP with a cost function $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}_{\geq 0}$ that quantifies safety violations. The objective is to maximize return while keeping the expected cumulative cost below a budget $\chi$:
$\max_{\pi}\;\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot\mid s)}[Q^{\pi}(s,a)]\quad\text{s.t.}\;\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot\mid s)}[Q_{c}^{\pi}(s,a)]\leq\chi,$ (CMDP)
Cost Budget. The budget $\chi$ specifies the maximum allowable expected cumulative cost and is typically treated as a human-selected threshold that reflects task-specific safety requirements [stooke2020responsive]. In this work, we assume a stricter setting by eliminating the cost budget, i.e., setting $\chi=0$, similar to [ganai2023iterative]. This corresponds to a hard-safety regime that aims to achieve minimal constraint violations during learning.
@@
$\pi^{*}=\arg\max_{\pi}\;\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot\mid s)}[Q^{\pi}(s,a)]\quad\text{s.t.}\;\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot\mid s)}[Q_{c}^{\pi}(s,a)]=0.$ (1)
@@ 4 Methodology
4.1 Split Architecture for Reward and Cost Optimization
@@
4.2 Action Modulation via Regulator Scaling
@@
where $\odot$ denotes element-wise multiplication; each component of the action vector is multiplied by a scaling factor between 0 (applied when high risk is predicted, resulting in large attenuation) and 1 (applied when the action is predicted to be safe, resulting in no attenuation), smoothly reducing potentially unsafe actions proportional to predicted risk. This element-wise modulation attenuates each component of the action based on its risk profile, reducing the magnitude of high-risk components. Unlike hard safety constraints that may override agent behavior, this approach preserves the agent’s exploration behavior and allows stable off-policy learning.
Assumptions. Our scaling mechanism assumes: (i) Monotonicity, where $Q_{c}(s,\rho\odot a)$ is expected to decrease as $\rho\to 0$, and (ii) Safety of Inaction, where zero-magnitude actions are safer than high-magnitude ones. These properties hold in many robotic domains where safety costs often scale with action magnitude, such as high torques causing actuator wear or large contact forces risking joint damage. By designing costs with this structure, scaling becomes an intuitive and effective tool for enforcing safety.
4.3 Learning Objectives and Updates
@@
$Q_{r}(s_{t},\tilde{a}_{t})\;\leftarrow\;$ $\;r(s_{t},\tilde{a}_{t})+\gamma\,\mathbb{E}_{\begin{subarray}{c}s_{t+1}\sim p(\cdot|s_{t},\tilde{a}_{t})a_{t+1}\sim\pi(\cdot|s_{t+1})\end{subarray}}\!\left[Q_{r}(s_{t+1},\tilde{a}_{t+1})\right],$ (3)
@@
ensuring that policy learning reflects the actual dynamics induced by the regulated action $\tilde{a}_{t}$. Our framework is algorithm-agnostic and can be integrated with any off-policy actor-critic method. For entropy-regularized algorithms such as SAC, the corresponding entropy term may be included in the actor objective. In our experiments, we demonstrate compatibility with both SAC and Twin Delayed DDPG (TD3)[fujimoto2018addressing].
Cost Learning. The cost critic is also trained on the scaled actions using a TD-style Bellman backup [sutton1998reinforcement]:
$Q_{c}(s_{t},\tilde{a}_{t})\;\leftarrow\;$ $\;c(s_{t},\tilde{a}_{t})+\gamma\,\mathbb{E}_{\begin{subarray}{c}s_{t+1}\sim p(\cdot|s_{t},\tilde{a}_{t})a_{t+1}\sim\pi(\cdot|s_{t+1})\end{subarray}}\!\left[Q_{c}(s_{t+1},\tilde{a}_{t+1})\right],$ (5)
@@
$\mathcal{L}_{\text{reg}}=\;\mathbb{E}_{s_{t}\sim\mathcal{D},\,a_{t}\sim\pi(\cdot|s_{t})}\Big[$ $\beta\cdot Q_{c}(s_{t},\tilde{a}_{t})-\lambda\cdot\log\rho_{\theta}(s_{t},a_{t},\hat{c}_{t})\,\Big].$ (6)
@@
Gradient Flow and Modularity. To ensure clean modularity, we detach the scaling weights $\rho_{\theta}(s_{t},a_{t},\hat{c}_{t})$ from the computational graph when updating both the reward and cost critics, preventing gradients from flowing through the regulator. Similarly, the actor receives no gradients from the regulator, learning purely from task returns. Moreover, the regulator is updated independently via its own objective, ensuring that reward maximization and safety modulation remain decoupled.
@@
Implementation Details. The RL agent follows its baseline implementation without modification. The regulator and twin cost critics are feedforward neural networks with two hidden layers of 256 units and ReLU activations. The regulator outputs element-wise scaling factors $\rho\in(0,1]^{d}$ via a sigmoid activation and is trained using the twin cost critics’ predictions.
5 Experiments
Figure 3: Performance comparison on the Safety Gymnasium locomotion environments. Each method is averaged over three independent runs; bold lines indicate the mean, and shaded areas show the standard deviation. Our methods (SAC-REG and TD3-REG) consistently achieve the best trade-off between return and cumulative constraint cost across all environments. Top: Episode return. Middle: Cumulative safety cost. Bottom: Return-to-Log-Cost ratio. Our methods outperform strong baselines, including PPO-PID [stooke2020responsive], SIMMER [sootla2022enhancing], SRCPO [kim2024spectral], RESPO [ganai2023iterative], and SEDITOR [yu2022towards]. SIMMER is omitted from the Swimmer plot as it consistently yields negative returns, moving opposite to the target velocity.